# Afford2Act: Affordance-Guided Automatic Keypoint Selection for Generalizable and Lightweight Robotic Manipulation

### v1

@@ Abstract
Vision-based robot learning often relies on dense image or point-cloud inputs, which are computationally heavy and entangle irrelevant background features. Existing keypoint-based approaches can focus on manipulation-centric features and be lightweight, but either depend on manual heuristics or task-coupled selection, limiting scalability and semantic understanding. To address this, we propose Afford2Act, an affordance-guided framework that distills a minimal set of semantic 2D keypoints from a text prompt and a single image. Afford2Act follows a three-stage pipeline: affordance filtering, category-level keypoint construction, and transformer-based policy learning with embedded gating to reason about the most relevant keypoints, yielding a compact 38-dimensional state policy that can be trained in 15 minutes, which performs well in real-time without proprioception or dense representations. Across diverse real-world manipulation tasks, Afford2Act consistently improves data efficiency, achieving an 82% success rate on unseen objects, novel categories, backgrounds, and distractors. Visualizations available at: https://afford2act.github.io/
@@ I Introduction
Learning effective robot manipulation policies hinges on overcoming a fundamental representation challenge: how to extract the manipulation-related features which are both compact and expressive? Dense visual inputs such as images [1], pointclouds [2], or semantic dense features [3, 4] provide rich details, but burden a policy with extraneous information. In contrast, 2D keypoint [5] offers a sparse state representation that highlights only the essential features of objects’ motions while still being representative enough to learn a good policy. This sparsity directly enables a lightweight policy with far fewer but important features than dense representation and support effective and real-time inference with less computational demand. However, the real challenge is in discovering these keypoints automatically and reliably: picking too few or irrelevant points can omit critical information, while relying on human insight for what points to use does not scale. In this paper, we present a method for identifying task-relevant object points that are both minimal (compact) and sufficient (expressive) for effective robot manipulation.
The problem of identifying keypoints has received a lot of attention recently. Some previous work utilizes manual selection [6] and prompts large Vision-Language Models (VLMs) to mark affordances on images [7, 3]. While these methods can provide useful cues, they depend on human guidance or external knowledge, making them labor-intensive and domain-specific. At the other end of the spectrum, some approaches forgo sparsity altogether in favor of dense feature grid-based representations. For instance, all previous work [8, 9, 10, 11] learn from dense grid of points in the scene or the object. While dense features are expressive, they are not compact and incur heavy computational demands. Furthermore, dense features may also entangle irrelevant dynamics (such as background features), which limits scalability and focus.
Figure 1: Top: policy trained with original prompts (“hold”, “stir”) executes the actions over time ($t$). Middle: at test time, the same policy generalizes to synonym prompts (“pick”, “mix”). Bottom: affordance-guided keypoints with learned importance (green; variable count; ring radius $\propto$ attention) remain stable under unseen object shapes, lighting conditions, dynamic distractors, and scene clutter—leading to successful rollouts.
Figure 2: Overview of the Afford2Act pipeline. We extract task-relevant object keypoints using affordance masks and DINO filtering, track them across frames, and embed them with a transformer and attention gating. The policy head predicts pose and gripper commands from these compact keypoint representations, enabling vision-based manipulation without requiring proprioception.
Recently, there has been work on automatically selecting a subset of keypoints from a dense grid [12, 13, 14] in an end-to-end manner. In particular, ATK [12] jointly optimizes a masking model for selecting a compact keypoint set and a task-specific control policy, yielding strong performance on challenging visuomotor tasks. However, there are two key limitations in ATK that prohibit it from generalizing: (i) Joint training couples the masking model to the task, so extending to a new task requires retraining both components (e.g., moving from mug picking from left to right); (ii) Existing evaluations emphasize robustness to positional variation and visual perturbations, whereas other generalization regimes, duplicate objects and cross-category transfer (e.g., from mugs to teapots), have received limited attention. In summary, prior works either require significant prior knowledge (manual keypoints), incur high computational overhead (dense grids), or struggle with stability and semantic generalization. This state of affairs motivates a core question for robot learning: How can a robot learn to identify the most relevant interaction points, the task hotspots, directly from visual inputs?
To address this challenge, we propose Afford2Act: Affordance-Guided Keypoints to Act, which uses text-prompted affordance cues to automatically select task-relevant keypoints, with no manual labels. This pipeline contains three stages: (i) Affordance Filtering (e.g., “hold” $\rightarrow$ mug handle), (ii) Keypoint Pool Construction from a small set of masked DINO anchors, and (iii) Keypoint-Based Policy Learning, where a transformer-based encoder followed by a gating mechanism learns the importance of each keypoint. Because mask support and geometry vary, the number of selected keypoints per object also varies; thus, correspondence is defined as a flexible set of affordance-region points. As visualized in Fig. 1, the policy typically attends to a small effective subset (Fig. 11) of keypoints. With a 38-dimensional input, policies remain robust under open-vocabulary prompts and scene changes across six real-world tasks. Evaluating across six real-world tasks, Afford2Act consistently outperforms keypoint-selection baselines and RGB/RGB-D policies, and stays robust and generalizable to synonym prompts, lighting, distractors, and clutter (Fig. 1).
@@
  1. 1.
    A unified Affordance-Guided Keypoint Distillation framework that extracts sparse, semantically meaningful 2D keypoints directly from text prompts.
  2. 2.
    An efficient, attention-based framework that trains a compact keypoint-guided policy in $\sim$15 minutes.
  3. 3.
    Extensive real-robot experiments in six tasks, demonstrating superior 82% success in unseen instances and robustness to open-vocabulary prompts (Fig. 1).
Figure 3: Real-world rollouts of Afford2Act on Pour (top) and Cut (bottom) tasks. In each sequence (left to right: approach $\rightarrow$ interaction $\rightarrow$ completion), green dots mark keypoints automatically selected on the affordance region (e.g., knife for “cut”, handle/spout for “pour”). The translucent ring size reflects the attention weight at each step. While keypoints remain fixed on task-relevant parts, the policy dynamically shifts its focus: for “cut”, the attention moves from the blade tip toward the handle; for “pour”, from the edge to the main handle. This adaptive attention enables robust, context-aware manipulation.
@@ II Related Work
Visual Representations for Policy Learning: Choosing the right visual representation is central to building robust manipulation policies. Classic approaches rely on raw sensory data which includes images, depth maps, or point clouds to provide a complete view of the environment [1, 2, 15]. While these representations are rich, they also saddle policies with high-dimensional inputs, requiring heavy data regimes and often resulting in policies that overfit to appearance rather than task. As a remedy, manipulation-centric and object-centric representations [3, 16, 17, 18, 19, 20, 21] filter out much of the visual clutter, focusing policy learning on the handful of features that matter most for action. However, these strategies often rely on hand-tuned detectors or dense input from images 3D reconstructions, limiting scalability and flexibility. Our method bypasses these constraints, extracting a compact, affordance-driven set of 2D keypoints that serve as a focused visual state, requiring neither dense geometry nor extensive supervision.
Keypoints and Automatic Anchor Discovery: Keypoints, when properly chosen, offer a natural low-dimensional representation that encodes an object’s essential affordances [6, 10, 11, 14]. Earlier work often assumed keypoints could be specified by experts, or required laborious labeling to maintain semantic consistency across objects [6, 22]. More recent efforts automate this selection, using object detector [10, 11] or learning which points are most predictive for the task at hand [12, 8]. For keypoint sampling from detected objects, the representations are still dense and lack semantic meaning. ATK, for instance, integrates keypoint selection with policy optimization, but can produce unstable and semantically ambiguous anchors. VLM-driven pipelines like MOKA [7] bring zero-shot power to keypoint discovery, but still rely on iterative visual prompting or domain knowledge. Our approach sidesteps both manual bias and unstable end-to-end discovery by using affordance cues as a reliable guide, automatically surfacing keypoints that are functionally meaningful and consistent, without relying on heavy annotation or black-box prompts.
@@ III Afford2Act
Overview. Given single-view demonstrations $\mathcal{D}=\{(I^{(d)}_{1:T_{d}},\,\mathbf{u}^{(d)}_{1:T_{d}},\,A^{(d)})\}_{d=1}^{D}$, where $I^{(d)}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images, $\mathbf{u}^{(d)}_{t}\in\mathbb{R}^{m}$ are actions (e.g., 6-DoF deltas and a gripper scalar), and $A^{(d)}$ is a text prompt, Afford2Act: (i) localizes the actionable region via a prompt-conditioned affordance mask (Section III-A), (ii) selects a small set of semantic keypoints on the reference frame (Section III-A), (iii) matches them within the mask on new instances to establish flexible correspondence ($\leq 15$) for various geometry and lightening (Section III-A), (iv) jointly tracks selected keypoints and encodes them into a transformer with a gating layer that learns phase-adaptive weights to encode a compact scene embedding for policy learning (Section III-B). The overall framework is shown in Fig. 2.
Formal Setup. Each demonstration $d$ is summarized by trajectories of $K$ 2D keypoints $\mathbf{P}^{(d)}=\{(x^{(d)}_{t,i},\,y^{(d)}_{t,i})\}_{t=1:T_{d},\;i=1:K}$, where indices $i$ refer to functional parts within the affordance region (e.g., “handle tip,” “blade edge”). This sparse, semantic representation lets the policy reason over roles rather than raw pixels.
@@ III-A Affordance-Guided Keypoint Representation
Affordance Localization: Given the first frame $I_{1}$ of a demonstration affordance prompt, we use One-Shot Open Affordance model  [28] $f_{\text{aff}}$ to produce a heatmap: $H=f_{\text{aff}}(I_{1},A)\in[0,1]^{H\times W}$. This heatmap indicates regions of the scene that likely correspond to the intended action. To focus only on the actionable surface, we apply a robust quantile threshold (along with mild smoothing) to obtain a binary mask: $M(x,y)=\mathbb{1}\big[H(x,y)\geq\tau_{q}(H)\big]$. This mask captures fine structures such as a knife blade or cup rim and robustly filters out clutter. Text-based prompts provide a strong semantic prior, allowing the same model to adapt affordance detection across novel tools or object categories with minimal (such as the examples in Fig. 3).
@@
Reference Keypoint Extraction: On a reference frame (in our case, the first frame of the first demonstration), $I_{1}^{r}$, we compute dense feature maps using a Vision Transformer backbone (ViT/DINO) [29]: $F_{r}=\phi(I^{r}_{1})\in\mathbb{R}^{H\times W\times D}$. We then restrict attention to the ROI defined by the affordance mask $M$ and cluster the features within that region to obtain $N$ reference anchor points:
$\mathcal{A}_{r}=\left\{(p_{i}^{r},\,\mathbf{f}_{i}^{r})\right\}_{i=1}^{N},\qquad p_{i}^{r}\in\Omega(M),\ \mathbf{f}_{i}^{r}=F_{r}(p_{i}^{r})$
@@
This step is fully automatic and scales to new categories, in contrast to manual annotation-based approaches [6, 7].
@@
Cross-instance Correspondence: Given a new demonstration, we extract features from its first frame, $F_{t}=\phi(I^{t}_{1})$. From each keypoint, $(p_{i}^{r},\mathbf{f}_{i}^{r})$ in the reference anchor, $\mathcal{A}_{r}$, we compute a cosine similarity map restricted to the affordance ROI in the new frame:
@@
$S_{i}(x,y)=\left\langle\frac{\mathbf{f}_{i}^{r}}{\|\mathbf{f}_{i}^{r}\|},\,\frac{F_{t}(x,y)}{\|F_{t}(x,y)\|}\right\rangle,\quad(x,y)\in\Omega(M_{t})$
We assign the keypoint with the highest similarity,
$p^{t}_{i}=\arg\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y),\quad c^{t}_{i}=\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y)$
@@
This matching ensures that the index $i$ refers to the same functional part (e.g., “handle tip”) across all demonstrations, even as object geometry and appearance vary. By constraining matches to the affordance mask, we avoid drifting to irrelevant but visually similar regions.
@@
Temporal Keypoint Tracking: Once all $K$ keypoints are initialized on the first frame, we track them across the demonstration using a transformer-based tracker such as CoTracker [30]: $\{p_{t,i}\}_{t=1:T,\,i=1:K}=g_{\text{trk}}(I_{1:T},\,\{p^{1}_{i}\}_{i=1}^{K})$
Joint tracking improves robustness to occlusion and viewpoint changes, leveraging correlations among keypoints. The output is a set of trajectories in a semantic space where each index $i$ always refers to the same actionable region.
By the end of this process, every demonstration is described by a trajectory of $K$ task-relevant, affordance-grounded keypoints. This sparse and semantically aligned representation serves as the input to our downstream policy (Section III-B), bridging the gap between the efficiency of keypoints and the semantic expressiveness of affordances, without requiring heavy 3D pipelines or manual keypoint curation.

### v2

@@ Abstract
Training vision-based manipulation policies that generalize beyond the training scenes remains challenging because dense visual observations entangle task-relevant cues with nuisance variation such as background, lighting, and instance-specific appearance. Keypoint representations offer a compact alternative, but automatically selecting task-relevant keypoints that transfer across objects and instructions is non-trivial. We present Afford2Act (Affordance-Guided Keypoints to Act), which uses language-conditioned affordance priors to localize where interaction should occur and extract a small set of semantic 2D keypoints that capture functionally relevant object parts. Since keypoint relevance can change over the course of an interaction, we use a lightweight attention-and-gating policy to prioritize keypoints for control. Across six real-robot manipulation tasks, Afford2Act achieves 82% success on held-out objects and is robust to clutter, distractors, and open-vocabulary prompt variations.
@@ I Introduction
Learning effective robot manipulation policies hinges on overcoming a fundamental representation challenge: how can we extract manipulation-related features that are both compact and expressive? Dense visual inputs such as images [4], point clouds [35], or semantic features [26, 30] provide rich detail, but may burden a policy with extraneous information [27]. In contrast, 2D keypoints [7, 20] offer a sparse input representation that highlights only the essential features of objects while still being representative enough to learn a good policy. This sparsity directly enables a lightweight policy with far fewer, but more important, features than dense representations and supports effective real-time inference with lower computational demand. However, the real challenge is discovering these keypoints automatically and reliably: picking too few or irrelevant points can omit critical information, while relying on human insight for point selection does not scale.
Automatic keypoint discovery for manipulation has recently gained momentum, with most methods falling into three paradigms. Motion-driven approaches identify keypoints as pixels (or point tracks) that move significantly when an object is manipulated, and then learn control from the resulting point tracks [2, 32, 34]. While effective when object motion is clean and informative, they can degrade in non-static scenes where background motion introduces spurious tracks, making it harder for the policy to remain reliable under dynamic distractors (e.g., a person operating in the same space as the manipulator). A second family of methods learns keypoints jointly with the policy, optimizing point selection end-to-end for task performance [36]. This can produce compact representations, but often couples the discovered points to the training data, making transfer to new objects or task intents less straightforward. Finally, a complementary direction seeks object-centric keypoints by extracting by grouping the semantic keypoints [29, 8]. These points can be helpful for consistent localization across frames and instances, but they are often agnostic to interaction: they describe where the object is and how it appears, rather than where the robot should act. Consequently, such keypoints may remain largely unchanged even when the intended use of the same object changes. In manipulation, however, relevance is inherently task-conditioned: the same tool supports multiple feasible interactions, and different instructions should highlight different regions (e.g., a brush can be grasped to lift/hold, or oriented and contacted to sweep; Fig. 4). This motivates an open question: can we automatically extract functionally grounded, task-conditioned keypoints that generalize across object shapes while remaining robust to distractors?
Figure 1: Top: Visualization of affordance-guided keypoint extraction with an example language prompt. Our pipeline reduces a $224\times 224$ image to 19 keypoints (38 dimensions). There are up to 15 keypoints on the object’s actionable region and 4 on the robot end-effector. The small sub-box in the top-right image is the affordance heatmap. Bottom: Examples of unseen instances to which our policy generalizes.
Figure 2: Overview of the Afford2Act pipeline. Step 1 takes single-view RGB demonstrations and language prompts as input. Step 2 performs affordance-guided keypoint distillation by localizing the actionable region and selecting reference anchor keypoints with DINO features. Step 3 establishes cross-instance correspondence to preserve semantic keypoint consistency across object instances. Step 4 tracks keypoints over time and learns a compact transformer-based policy with gating to prioritize task-relevant tokens when predicting end-effector motion and gripper commands.
To address this challenge, we turn to affordances [26, 18]. A natural way to make feasible interactions explicit is to link an instruction (e.g., hold, cut, pour) to actionable regions in the scene. Building on this intuition, we introduce AFFORD2ACT: Affordance-Guided Keypoints to Act. Our approach first filters candidate points to instruction-relevant regions, then preserves functional keypoint identity across object instances through correspondence, and finally adapts keypoint importance during control with gating that dynamically re-weights keypoints over time, focusing on different interaction points at different stages of execution. Together, these components produce a compact state tied to functionally meaningful object parts rather than appearance, enabling efficient policy learning from a single external RGB view. Across six real-world manipulation tasks, this representation remains robust to novel objects, clutter, and open-vocabulary prompts.
@@
  • A language-conditioned affordance-guided keypoint extraction pipeline that selects sparse 2D keypoints from interaction-relevant regions without manual keypoint specification (Sec. III-A).
  • A correspondence-and-gating policy formulation that preserves keypoint consistency across object instances while adapting keypoint relevance over the course of manipulation (Sec. III-B).
  • Extensive real-robot evaluation on six tasks, showing robust transfer to held-out objects, scene variations, and prompt paraphrases with a compact keypoint state (Secs. IV, IV-E; Fig. 1).
Figure 3: Real-world rollouts of Afford2Act on unseen instances. Rows show Pour (top), Cut (middle), and Stir (bottom), with each sequence progressing left-to-right from approach $\rightarrow$ interaction $\rightarrow$ completion. Green dots denote affordance-selected keypoints, and translucent ring size indicates attention weight over time. Keypoints remain anchored to task-relevant parts, while attention shifts with phase (e.g., cut: blade tip $\rightarrow$ handle; pour: edge/spout $\rightarrow$ main handle; stir: handle $\rightarrow$ bowl), enabling robust, context-aware manipulation.
@@ II Related Work
Visual Representations for Policy Learning: Choosing the right visual representation is central to building robust manipulation policies. Classic approaches rely on raw sensory inputs such as images, depth maps, or point clouds to provide a complete view of the environment [35]. While these representations are rich, they also saddle policies with high-dimensional inputs, requiring large data regimes and often resulting in policies that overfit to appearance rather than task. As a remedy, manipulation-centric and object-centric representations [26, 13, 9, 24, 33, 25, 12] filter out much of the visual clutter, focusing policy learning on the handful of features that matter most for action. However, these strategies often rely on hand-tuned detectors or dense inputs from images or 3D reconstructions, limiting scalability and flexibility. Our method bypasses these constraints by extracting a compact, affordance-driven set of 2D keypoints that serves as a focused visual state, requiring neither dense geometry nor extensive supervision.
Keypoints and Automatic Anchor Discovery: Keypoints, when properly chosen, offer a natural low-dimensional representation that encodes an object’s essential affordances [17, 32, 34, 29]. Earlier work often assumed keypoints could be specified by experts or required laborious labeling to maintain semantic consistency across objects [17, 20]. More recent efforts automate this selection, using object detectors [32, 34] or learning which points are most predictive for the task at hand [36, 31]. When keypoints are sampled from detected objects, the resulting representations are still dense and often lack semantic meaning. ATK, for instance, integrates keypoint selection with policy optimization, but can produce unstable and semantically ambiguous anchors. VLM-driven pipelines like MOKA [19] bring zero-shot power to keypoint discovery, but still rely on iterative visual prompting or domain knowledge. Our approach sidesteps both manual bias and unstable end-to-end discovery by using affordance cues as a reliable guide, automatically surfacing keypoints that are functionally meaningful and consistent, without relying on heavy annotation or black-box prompts.
@@ III Afford2Act
Overview. Given single-view demonstrations $\mathcal{D}=\{(I^{(d)}_{1:T_{d}},\,\mathbf{u}^{(d)}_{1:T_{d}},\,A^{(d)})\}_{d=1}^{D}$, where $I^{(d)}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images, $\mathbf{u}^{(d)}_{t}\in\mathbb{R}^{m}$ are actions (e.g., 6-DoF deltas and a gripper scalar), $T_{d}$ is the timestep at the end of the demonstration, and $A^{(d)}$ is an affordance text prompt, Afford2Act: (i) localizes the actionable region via a prompt-conditioned affordance mask (Section III-A), (ii) selects a small set of semantic keypoints on the reference frame (Section III-A), (iii) matches them within the mask on new instances to establish flexible correspondence with up to 15 keypoints across variations in geometry and lighting (Section III-A), and (iv) jointly tracks selected keypoints and encodes them into a transformer with a gating layer that learns phase-adaptive weights to produce a compact scene embedding for policy learning (Section III-B). The overall framework is shown in Fig. 2.
Setup. Each demonstration $d$ is summarized by trajectories of $K$ 2D keypoints, represented as $\mathbf{P}^{(d)}=\{(x^{(d)}_{t,i},\,y^{(d)}_{t,i})\}_{t=1:T_{d},\;i=1:K}$, where indices $i$ refer to functional parts within the affordance region (e.g., “handle tip,” “blade edge”). This sparse semantic representation lets the policy reason over roles rather than raw pixels.
@@ III-A Affordance-Guided Keypoint Representation
Affordance Localization: Given the first frame $I_{1}$ of a demonstration and an affordance prompt $A$, we use a pre-trained One-Shot Open Affordance model [18] $f_{\text{aff}}$ to produce a heatmap: $H=f_{\text{aff}}(I_{1},A)\in[0,1]^{H\times W}$. This heatmap indicates regions of the scene that correspond to the intended action. To focus only on the actionable surface, we apply a robust quantile threshold (along with mild smoothing) to obtain a binary mask: $M(x,y)=\mathbb{1}\big[H(x,y)\geq\tau_{q}(H)\big]$, where $\tau_{q}(H)$ is the threshold value. This mask captures fine structures such as a knife blade or cup rim and robustly filters out clutter.
@@
Reference Keypoint Extraction: On a reference frame (the first frame of the first demonstration), $I_{1}^{r}$, we compute dense feature maps using a DINO backbone ($\phi$[22]: $F_{r}=\phi(I^{r}_{1})\in\mathbb{R}^{H\times W\times D}$. We use DINO features because they provide a robust, part-sensitive similarity metric that enables more reliable cross-instance correspondence. We then restrict attention to the ROI  [^1] ROI stands for Region of Interest, i.e., the cropped region of the affordance mask. defined by the affordance mask $M$ and cluster the features within that region to obtain $K$ reference anchor points in a reference anchor set $\mathcal{A}_{r}$:
$\mathcal{A}_{r}=\left\{(p_{i}^{r},\,\mathbf{f}_{i}^{r})\right\}_{i=1}^{K},\qquad p_{i}^{r}\in\Omega(M),\ \mathbf{f}_{i}^{r}=F_{r}(p_{i}^{r})$
@@
where $\Omega(M)$ is the set of pixel coordinates covered by the mask $M$. This step is fully automatic and scales to new categories, in contrast to manual annotation-based approaches [17, 19].
@@
Cross-instance Correspondence: Given a new demonstration, we compute a dense feature map from its first frame, $F_{t}=\phi(I^{t}_{1})$. For each reference anchor keypoint $(p_{i}^{r},\mathbf{f}_{i}^{r})\in\mathcal{A}_{r}$, we match it in the new frame by evaluating cosine similarity between the reference descriptor and the per-pixel features, restricted to the affordance ROI:
@@
$S_{i}(x,y)=\left\langle\frac{\mathbf{f}_{i}^{r}}{\|\mathbf{f}_{i}^{r}\|},\,\frac{F_{t}(x,y)}{\|F_{t}(x,y)\|}\right\rangle,\quad(x,y)\in\Omega(M_{t}).$
We then select the best-matching location and its confidence as:
$p^{t}_{i}=\arg\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y),\quad c^{t}_{i}=\max_{(x,y)\in\Omega(M_{t})}S_{i}(x,y).$
@@
This matching ensures that the index $i$ refers to the same functional part (e.g., “handle tip”) across demonstrations, even as object geometry and appearance vary. By constraining matches to the affordance mask, we reduce spurious matches to visually similar but task-irrelevant regions.
@@
Temporal Keypoint Tracking: After initializing all $K$ keypoints in the first frame, we track them over the demonstration using a transformer-based point tracker (e.g., CoTracker [14]). This produces a temporally ordered set of keypoint locations,
$\{(p_{1,1},\ldots,p_{1,K}),\;(p_{2,1},\ldots,p_{2,K}),\;\ldots,\;(p_{T_{d},1},\ldots,p_{T_{d},K})\},$
where $p_{t,i}$ denotes the 2D location of keypoint $i$ at time $t$ and $T_{d}$ is the demonstration length. Joint keypoint tracking with Co-Tracker improves robustness to occlusion (e.g., when the gripper blocks the view of the handle during grasping) by exploiting correlations among keypoints, and preserves index consistency so that each $i$ continues to refer to the same actionable region over time. After this step, each demonstration is represented as a trajectory of $K$ task-relevant, instruction-conditioned keypoints. We use this sparse, temporally consistent representation as input to the downstream policy (Section III-B), without requiring dense 3D reconstruction or manual keypoint specification.